OpenStudy (anonymous):

Solve using the quadratic formula. Approximate answers to the nearest tenth.2x^2-3x-8=0

OpenStudy (mathstudent55):

This is the quadratic formula: \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) It is used to solve an equation of the form \(ax^2 + bx + c = 0\) Your equation is in that form with a = 2, b = -3, and c = -8. Just use those numbers for a, b, and c in the formula above and find what x is equal to.